# CoRa: A Collision-Resistant LoRa Symbol Detector of Low Complexity

### v1

@@ IV Evaluation
To foster a fair comparison, we conduct the analysis on the two datasets provisioned by the related work. Both datasets contain real-world LoRa signals of fixed transmitter devices. The first dataset (from CIC) includes four scenarios with varying SNR conditions and fixed PHY parameters. The second dataset (from TnB) comprises three scenarios with varying SNR conditions and varying PHY parameters. The characteristics of these datasets, as well as the transmission parameters, are summarized in Table I.
@@
Spreading Factor (SF) 8 8, 10
Bandwidth (B) 250 kHz 125 kHz
Coding Rate (CR) 4/5 4/5, 4/6, 4/7, 4/8
Payload Size 12 B 12 B
Sampling Rate 2 MHz 1 MHz
SNR (SF8) 30-42 dB 30-42 dB 10-30 dB -17-5 dB -5-17 dB -10-10 dB -8-14 dB
SNR (SF10) - -9-23 dB -17-12 dB -13-12 dB
Number of TX Devices 20 20
TX Rate 5-100 pkt/s 20, 25 pkt/s
Capture Time 60 s 30 s
@@ IV-A Collision Resolution
Figure 6: Throughput as a function of aggregated transmission rate across different scenarios in the CIC dataset. Devices transmit using a spreading factor of 8, a bandwidth of 250 KHz and a coding rate of 4/5. The shaded areas around the curves indicate the error margins.
Figure 6 depicts the throughput of the baseline, CIC, TnB, and the CoRa demodulator using the CIC dataset. The CoRa approach demonstrates superior throughput in all scenarios, achieving up to 29% improvement over TnB in high SNR conditions without Line-of-Sight (see Figure 6, D2 at 95 pkt/s) and up to a 6% improvement in the Sub-Noise SNR scenario (see Figure 6, D4 at 100 pkt/s). Given that our demodulator utilizes the synchronization and decoding blocks of TnB, the observed performance difference is attributed to the superior efficacy of our demodulator over the Thrive demodulator utilized in TnB. Consequently, our solutions exhibits up to a 178% improvement over CIC and up to 11.53 times better throughput than the baseline in high SNR conditions (see Figure 6, left, 100 pkt/s).
@@
Figure 7: Throughput as a function of coding rate across different scenarios in the TnB dataset. Devices transmit using a spreading factor of 8 and 10, a bandwidth of 125 KHz and transmission rate of 25 packets per second. Almost all captures use this rate, except for those with a spreading factor of 10 and coding rates of 4/5 and 4/8, which use a transmission rate of 20 packets per second.
@@ IV-B Resistance to Multipath Fading and Doppler Effect
Figure 8: Packet Reception Ratio as a function of coding rate across different scenarios in the simulated dataset, for the LTE Extended Typical Urban (ETU) channel and a Flat channel. Devices transmit using a spreading factor of 8 and 10, a bandwidth of 125 KHz, and transmission rate of 15 pkt/s.
@@
Figure 8 presents the packet reception ratio for the evaluated demodulators. In the SF8 scenario under the ETU channel conditions (top-left), our demodulator outperforms the current state-of-the art solutions, despite the strong strong channels fluctuations. Specifically, our demodulator improves the packet reception ratio over TnB by approximately 12.3% with a coding rate of 4/5 and by about 5% with a coding rate of 4/8. In the SF10 scenario (top-right in the figure), both the CoRa and the TnB demodulators perform similarly, with less than one percent difference.
@@ IV-C Performance overhead
Figure 9: Execution time per symbol on the target machine (Quad-Core Intel Core i7-8650U) for the standard LoRa demodulator (Baseline) and the CoRa demodulator, across two spreading factors (8 and 10). The contribution of each stage to the total demodulation time is illustrated.
@@ V Discussion
Moreover, the independence of symbol boundary information makes our symbol detector versatile in various scenarios. One such scenario is Frequency Hop Spreading Spectrum (FHSS), a feature supported by common LoRa transceivers. Another scenario involves accurately detecting symbols even if the frame synchronization mechanism fails to detect colliding frames. In case of FHSS, when the frequency hopping pattern is unknown – such as when independent LoRa networks transmit data simultaneously – it becomes impossible to predict potential symbol collisions since symbols may appear on different channels. This limitation renders certain algorithms ineffective, such as TnB or CoLoRa, which are designed to match peaks to transmitting devices.
@@
Our demodulator exhibits significantly lower computational complexity than the CIC symbol detector, which requires $(1+2K)$ Fast Fourier Transforms (FFTs) per colliding symbol, where $K$ is the number of detected colliding frames. Hence, the complexity of CIC increases polynomially with the number of collisions, whereas our solution remains constant. Thrive (TnB) involves peak detection, curve smoothing using Robust Locally Weighted Scatterplot Smoothing (rlowess), and matching cost evaluations. While peak detection and cost evaluations are simple, rlowess involves weighted linear regressions, which are computationally more demanding than FFTs. Although the exact computational performance of Thrive and CoRa depend on the implementation, rlowess is sequential and limits parallelization. In contrast, the FFTs in our demodulator are computationally efficient and inherently can run in parallel, as many FFT implementations leverage the Radix algorithm, which is well-known for its capability to be efficiently parallelized.
@@ VI Conclusions and outlook
Results indicated that our approach yields outcomes that are either superior or at least comparable with those of the state-of-the-art solutions, even under conditions of sub-noise Signal-to-Noise Ratio (SNR). The performance advantage of CoRa is particularly evident in scenarios characterized by extreme collisions. Notably, our solution maintains a constant overhead, specifically three times the demodulation times of the baseline LoRa demodulator. This suggest the potential feasibility of the implementation on real hardware.
@@
All implementations, simulations, and datasets used in this study are publicly available for review and reproduction at https://zenodo.org/records/14515243

### v2

@@ IV Evaluation
To foster a fair comparison, we conduct the analysis on the two datasets provisioned by the related work. Both datasets contain real-world LoRa signals of fixed transmitter devices. The first dataset (from CIC) includes four scenarios with varying SNR conditions and fixed PHY parameters. The second dataset (from TnB) comprises three scenarios with varying SNR conditions and varying PHY parameters. The characteristics of these datasets, as well as the transmission parameters, are summarized in Table I. All captures involve 20 transmitter devices, each sending a 12-byte payload.
@@
Spreading Factor 8 8, 10
Bandwidth [kHz] 250 125
Coding Rate 4/5 4/5, 4/6, 4/7, 4/8
Sampling Rate [MHz] 2 1
SNR (SF8) [dB] 30 to 42 30 to 42 10 to 30 -17 to 5 -5 to 17 -10 to 10 -8 to 14
SNR (SF10) [dB] - -9 to 23 -17 to 12 -13 to 12
Deployment area [m2] $15\times 10$ $100\times 60$ $170\times 100$ $1200\times 1600$ $100\times 120$ $400\times 240$ $220\times 200$
TX Rate [pkt/s] 5 to 100 20, 25
Capture Time [s] 60 30
@@ IV-A Collision Resolution
Figure 6: Throughput vs aggregated transmission rate in the CIC dataset. The shaded areas around the curves indicate the error margins.
Figure 6 depicts the throughput of the baseline, CIC, TnB, and the CoRa demodulator using the CIC dataset. The CoRa approach demonstrates superior throughput in all scenarios, achieving up to 29% improvement over TnB in high SNR conditions without Line-of-Sight (see Figure 6, D2 at 95 pkt/s) and up to a 6% improvement in the Sub-Noise SNR scenario (see Figure 6, D4 at 100 pkt/s). Given that our demodulator utilizes the synchronization and decoding blocks of TnB, the observed performance difference is attributed to the superior efficacy of our demodulator over the Thrive demodulator utilized in TnB. Correspondingly, our solutions exhibits up to a 178% improvement over CIC and delivers up to 11.53 times higher throughput than the baseline in high SNR conditions (see Figure 6, left, 100 pkt/s).
@@
Figure 7: Throughput vs coding rate in the TnB dataset. Devices with spreading factor 10 and coding rates 4/5 and 4/8 transmit 20 pkt/s, while devices in al other scenarios transmit 25 pkt/s.
@@ IV-B Resistance to Multipath Fading and Doppler Effect
Figure 8: Packet Reception Ratio vs. coding rate in the simulated dataset for LTE Extended Typical Urban (ETU) and Flat channels. Devices transmit at 15 pkt/s with SF8 (SNR: $[-5,15]$ dB) and SF10 (SNR: $[-10,10]$ dB), using a 125 kHz bandwidth.
@@
Figure 8 presents the packet reception ratio for the evaluated demodulators. In the SF8 scenario under the ETU channel conditions (top-left), our demodulator outperforms the current state-of-the art solutions, despite the strong channel fluctuations. Specifically, our demodulator improves the packet reception ratio over TnB by approximately 12.3% with a coding rate of 4/5 and by about 5% with a coding rate of 4/8. In the SF10 scenario (top-right in the figure), both the CoRa and the TnB demodulators perform similarly, with less than one percent difference.
@@ IV-C Performance Overhead
Figure 9: Execution time per symbol for standard LoRa (Baseline) and CoRa demodulators at SF8 and SF10, showing stage-wise contribution to total demodulation time.
@@ V Discussion
Moreover, the independence from symbol boundary information enhances the versatility of our symbol detector across various scenarios. One such scenario is Frequency Hop Spreading Spectrum (FHSS), where the LoRa transceiver transmit packet fragments across multiple channels. This feature is standard in LoRa devices and should not be confused with LR-FHSS, which employs GMSK modulation and narrowband signals. Another scenario involves accurately detecting symbols even if the frame synchronization mechanism fails to detect colliding frames. In case of FHSS, when the frequency hopping pattern is unknown – such as when independent LoRa networks transmit data simultaneously – it becomes impossible to predict potential symbol collisions since symbols may appear on different channels. This limitation renders certain algorithms ineffective, such as TnB or CoLoRa, which are designed to match peaks to transmitting devices.
@@
Our demodulator exhibits significantly lower computational complexity than the CIC symbol detector, which requires $(1+2K)$ Fast Fourier Transforms (FFTs) per colliding symbol, where $K$ is the number of detected colliding frames. Hence, the complexity of CIC increases polynomially with the number of collisions, whereas our solution remains constant.
The Thrive (TnB) algorithm involves three main computational steps: FFT calculation ($O(NlogN)$), peak detection ($O(N)$), and peak cost evaluations ($O(N)$). These peak cost evaluations are performed up to $2M^{2}$ times per symbol, where $M$ denotes the number of colliding packets within the symbol. Since $M\ll N$, where N denotes the number of samples per symbol, the overall complexity of Thrive simplifies to $O(NlogN)$, aligning with our solution, where the FFT step dominates the total computational cost. Still, the runtime of TnB cost evaluations scales quadratically, leading to unbounded execution time. In our solution, the runtime remains constant.
@@ VI Conclusions and outlook
Results indicated that our approach yields outcomes that are either superior or at least comparable with those of the state-of-the-art solutions, even under conditions of sub-noise Signal-to-Noise Ratio (SNR). The performance advantage of CoRa is particularly evident in scenarios characterized by extreme collisions. Notably, our solution maintains a constant overhead, specifically three times the demodulation times of the baseline LoRa demodulator. This suggests the potential feasibility of the implementation on real hardware.
@@
The authors have provided public access to their code and data at https://doi.org/10.5281/zenodo.14582713